OpenStudy (anonymous):

Rupert's backyard is 10 m by 5 m. The area of the yard is: 50 m 50 m^2 15 m 15 m^2

OpenStudy (anonymous):

I think it is b

OpenStudy (jhannybean):

A= L*W L = 10 m W = 5 m A = (10)(5) = 50 m^2

OpenStudy (anonymous):

a=l*b square units

OpenStudy (mathstudent55):

If you keep the units in the formula for the area, you'll get them right. Assuming this is a rectangular backyard, the area id given by the length times the width: A = L * W Since we are told L = 10 m and W = 5 m, plug them into the formula with the units: A = 10 m * 5 m = 10 * 5 * m * m = 50 m^2